Given the functions: ATC= x-6+(30/x) and AR =10+(15/x), where x represents output, find TC, TR, and the output level when net revenue is zero.



1
Expert's answer
2021-02-17T12:10:14-0500

"AR =10+(15\/x)"

"ATC= x-6+(30\/x)"

"ATC=TC\/X"

"TC=ATC.X"

"TC=(X-6+30\/X)X"

"TC=X.X-6X+30"


"AV=TR\/X"

"TR=AV.X"

"TR=(10+15\/X)X"

"TR=10X+15"


"Net revenue=TR-TC"

"10X+15-(X.X-6X+30)"

"X.X-16X+15=0"

"X(X-15)-1(X-15)=0"

"X-1=0 X=1"

"X-15=0 X=15"

The output level when net output is zero is 15 or 1